# Advancing Glitch Classification in Gravity Spy: Multi-view Fusion with Attention-based Machine Learning for Advanced LIGO’s Fourth Observing Run

### v1

@@ Abstract
The first successful detection of gravitational waves by ground-based observatories, such as the Laser Interferometer Gravitational-Wave Observatory (LIGO), marked a revolutionary breakthrough in our comprehension of the Universe. However, due to the unprecedented sensitivity required to make such observations, gravitational-wave detectors also capture disruptive noise sources called glitches, potentially masking or appearing as gravitational-wave signals themselves. To address this problem, a community-science project, Gravity Spy, incorporates human insight and machine learning to classify glitches in LIGO data. The machine learning classifier, integrated into the project since 2017, has evolved over time to accommodate increasing numbers of glitch classes. Despite its success, limitations have arisen in the ongoing LIGO fourth observing run (O4) due to its architecture’s simplicity, which led to poor generalization and inability to handle multi-time window inputs effectively. We propose an advanced classifier for O4 glitches. Our contributions include evaluating fusion strategies for multi-time window inputs, using label smoothing to counter noisy labels, and enhancing interpretability through attention module-generated weights. This development seeks to enhance glitch classification, aiding in the ongoing exploration of gravitational-wave phenomena.
keywords:
LIGO, Glitch Classification, Gravity Spy, O4, Machine Learning, Attention
journal: Information Science
@@ 1 Introduction
The first discovery of gravitational waves, a pivotal element in Einstein’s theory of general relativity einstein1916approximative , has opened up an entirely new window in the cosmos. The Laser Interferometer Gravitational-Wave Observatory (LIGO) aasi2015advanced achieved groundbreaking success in detecting these ripples in spacetime for the first time in September 2015 abbott2016observation , fundamentally transforming our understanding of the Universe. Later on, Virgo acernese32advanced made its first detection of gravitational waves in August 2017 PhysRevLett.119.141101 ; abbott2019gwtc and KAGRA initiated its gravitational waves observations in April 2020 kagra2019kagra ; ligo2022first . However, achieving these observations demands exceptionally sensitive and intricate detectors to be able to measure minuscule fluctuations in spacetime abbott2020guide . This heightened sensitivity, in turn, leads to the detection of diverse sources of noise, with the potential to obscure or mimic authentic gravitational-wave signals abbott2016characterization ; davis2021ligo . Among these noises, bursts of non-Gaussian noise caused by environmental or instrumental factors, known as glitches, are particularly disruptive to measuring gravitational waves. The origins of many glitches remain largely unknown davis2022detector ; nuttall2018characterizing . The process of characterizing and eliminating these glitches is essential to enhancing the quality of the detection system, thereby increasing the likelihood of analyzing gravitational waves.
@@
In pursuit of this objective, Gravity Spy zevin2017gravity ; zevin2023gravity , a community-science project, has been launched to leverage the power of both volunteers and machine learning to classify glitches in the LIGO time-series data. On one front, the project engages human expertise in identifying and categorizing various glitches collected by LIGO. These labeled glitches then serve as training data to refine machine learning classifiers, enhancing the accuracy and efficiency of glitch identification. Conversely, the classifiers provide guidance to volunteers, assisting them in identifying existing or new potential glitch classes coughlin2019classifying . Gravity Spy is implemented on the Zooniverse platform fortson2018optimizing on the community-science side, and it has demonstrated significant success by attracting over 30 thousand volunteers making more than 7 million glitch classifications.111gravityspy.org The outputs are actively used by LIGO. One critical component contributing to this significant achievement is the integration of the machine learning classifier.
@@
The machine learning classifier has been seamlessly integrated into the Gravity Spy pipeline for over five years, largely maintaining its original architecture. Periodically, fine-tuning of its final layer has been performed to enhance its adaptability to diversely predicted glitch classes dodge2020fine . The input for the classifier remains unchanged, featuring a structure that juxtaposes four time window spectrograms (0.5 s, 1.0 s, 2.0 s and 4.0 s) of a single glitch into one merged image. The first version of the machine learning classifier was introduced in the initial Gravity Spy paper zevin2017gravity and integrated into the Gravity Spy pipeline after the completion of the first observing run (O1). Designed for a 20-class classification setup, this classifier consisted of two convolutional layers and two fully-connected layers to extract image features. Each convolutional layer was followed by a max pooling layer to reduce the dimensionality of the features.
@@
As the project progressed, during the second observing run (O2), Gravity Spy volunteers identified new glitch classes and two of them, i.e., 1080 Lines and 1400 Ripples, were added into the classifier. Therefore, the classifier was optimized by incorporating two additional convolutional layers and max pooling layers before the existing fully-connected layers, which facilitated the extraction of more discriminative features and adapted the classifier to a 22-class classification task bahaadini2018machine .
@@
During the third observing run (O3), two new glitch classes, known as Low Frequency Blip and Fast Scattering, were identified through the collaborative efforts of volunteers and LIGO experts soni2021discovering . Alongside this, the decision was made to remove the None of the Above class from the training dataset. The None of the Above class was initially aimed at empowering our volunteers to identify new glitch classes. However, its inclusion confuses the model, especially given morphological similarities with other glitches. As a result, removing this class allows the identification of new glitches more effectively with low predicted confidence soni2021discovering ; glanzer2023data . This led to further retraining of the classifier, enabling it to effectively accommodate a 23-class classification task soni2021discovering .
@@
Currently, the fourth observing run (O4) of the LIGO detector network is in process abbott2020prospects . With the anticipation of discovering new glitch classes during O4 and subsequent runs, the current Gravity Spy classifier faces major limitations alvarez2023gspynettree ; jarov2023new . First, in the glitch classification task, unlike other image processing tasks, location plays a crucial role. For example, distinguishing between a line at the top and a line at the bottom of the image is essential as they represent different glitches at different frequencies. However, the classifier’s relatively straightforward architecture, comprised mainly of a combination of convolutional and fully-connected layers, hinders its capacity to capture unique features in individual glitch classes simonyan2014very . Second, due to its shallow layer structure, the classifier requires significant resizing of input image dimensions, resulting in the loss of valuable information. Moreover, the classifier uses a single model to extract features from merged image inputs across different time windows, which further constrains the ability to uncover cross-time window correlations huang2020fusion . Recognizing these challenges, it becomes apparent that the demand to analyze more data and diverse glitch types exceeds the current model’s capabilities.
@@
Additionally, the current classifier grapples with a confidence overfitting issue, displaying excessive certainty in glitch predictions even when errors are present. This tendency can potentially mislead both volunteers and LIGO experts, posing a risk to the efficiency and reliability of glitch characterization. Therefore, there is a compelling need for an advanced classifier architecture with robust generalization capabilities in the context of O4.
@@
In this study, we develop a novel machine learning classifier for glitch classification in the ongoing O4 run to address these challenges. The proposed classifier employs an attention-based multi-view fusion strategy to capture cross-time window correlations across the four time windows. The incorporation of a regularizer and label smoothing techniques into the loss function improves generalization and mitigates the confidence overfitting issue. The contributions of this paper can be summarized as follows: 1) We comprehensively compare three fusion strategies (early fusion, late fusion, and intermediate fusion) in scenarios with multiple time windows as inputs. 2) We apply label smoothing to mitigate the impact of noisy training labels. 3) We introduce attention modules in the classifier, enhancing transparency in the glitch decision-making process by identifying specific window time that predominantly influences the glitch classification decision — an exploration undertaken for the first time.
@@ 2 Materials and Methods
In this section, we provide an overview of the Gravity Spy dataset, covering its data generation process, image preprocessing, and the methodology for data splitting during both training and testing phases. Following this, we delve into the multi-view fusion strategy, discussing its early, intermediate, and late fusion components. Last, we propose a novel classifier, outlining the model architecture, attention modules, and the application of a label smoothing technique.
@@ 2.1 Gravity Spy Dataset
The Gravity Spy dataset comprises time-frequency spectrograms of glitches. The glitches are initially identified by the Omicron trigger pipeline robinet2016omicron . To effectively capture glitches that pose significant challenges to gravitational-wave searches, this pipeline employs strict criteria, identifying only noise events with a signal-to-noise ratio exceeding 7.5 and a peak frequency falling within the range of 10 Hz to 2048 Hz. Subsequently, the Q-transform process chatterji2004multiresolution ; chatterji2005search is employed to convert the LIGO strain data into time-frequency spectrograms, known as Omega scans. This involves segmenting the strain time series into shorter intervals and applying the discrete Fourier transform to extract the frequency content. Gravity Spy generates four spectrograms for each glitch triggered, using distinct time windows: 0.5 s, 1.0 s, 2.0 s, and 4.0 s. The incorporation of multiple time windows serves a crucial purpose, allowing both volunteers and machine learning models to analyze the morphologies of glitches occurring at different characteristic timescales. For instance, this enables the examination of short (i.e., Blip) and long duration (i.e., Whistle) noise events, as depicted in Fig. 1.
@@
Figure 1: Spectrogram examples of two glitches with four time windows (0.5 s, 1.0 s, 2.0 s and 4.0 s). (a) Blip, characterized by its short duration. (b) Whistle, characterized by its long duration. These examples illustrate how different time windows enable the analysis of glitches occurring at various temporal scales, providing valuable insights into the morphological characteristics of each glitch type.
@@
The training and testing datasets employed in this study are sourced from Zevin et al. zevin2022gravity , focusing exclusively on data from the Hanford and Livingston detectors during the third observing run (O3) of Advanced LIGO. This classification task has in total 23 glitch classes. To ensure consistency, the names associated with each class and the typical morphology of glitches belonging to each class were determined through a collaborative effort between LIGO scientists and citizen scientists. Examples representing each of the 23 glitch classes are illustrated in Fig. 2 and more details are provided in the O1-O3 data release paper glanzer2023data .
@@
Figure 2: Spectrogram examples of each of the 23 glitch classes within the Gravity Spy dataset. All glitches share the same axes and scale bar with Violin Mode. These examples provide one possible morphological representation for each class, while each class can exhibit various shapes and patterns glanzer2023data .
@@
For the purposes of this study, we consider the ground truth for glitches in our dataset to be established through a collaborative effort involving volunteers and the previously implemented machine learning classifier zevin2017gravity . However, this ground truth may exhibit a potential noisy-label issue, that is a few glitches were misclassified by the classifier and some volunteers. To strengthen the confidence in the ground truth used for our new classification model, we adopt a filtering criterion to create the training and testing dataset. Specifically, we only use glitches with consensus confidence scores from both machine learning and volunteers greater than 0.9. As a consequence, our new model is trained on glitches with relatively reliable labels, effectively reducing confusion during the training process and shielding the model from potentially mislabeled ground truth. Improved training sets for O4 data will be the focus of future studies.
@@
Following the filtering criteria, the Gravity Spy dataset is divided into three distinct sets. The training set consists of 8439 glitches and the validation set consists of 1687 glitches. Additionally, a separate set of 3538 glitches is reserved for testing purposes. Each set is carefully curated to ensure no overlapping glitch subjects, thereby preventing data leakage between those three phases. A detailed breakdown of the data splitting is presented in Table 1.
@@
Table 1: The distribution of each class in the Gravity Spy dataset across the training, validation, and testing sets.
Class Train Validation Testing Class Train Validation Testing 1080Lines 184 49 225 Low Frequency Lines 381 91 225 1400Ripples 257 64 224 No Glitch 311 78 225 Air Compressor 293 57 225 Paired Doves 124 31 146 Blip 312 84 218 Power Line 354 83 224 Low Frequency Blip 337 103 222 Repeating Blips 411 90 224 Chirp 38 10 17 Scattered Light 521 128 224 Extremely Loud 327 85 224 Scratchy 178 36 225 Fast Scattering 515 128 224 Tomte 408 90 225 Helix 46 10 71 Violin Mode 210 56 225 Koi Fish 338 93 224 Wandering Line 25 8 25 Light Modulation 127 28 159 Whistle 462 129 224 Low Frequency Burst 378 106 223
@@
The glitches from all three cohorts undergo the same pre-processing steps. Initially, each spectrogram’s original dimension size is $600\times 800\times 3$. We first crop the spectrogram with a bounding box defined as [top-left vertical:60, top-left horizontal:100, height:580, width:675] to retain only the intensity region of interest and eliminate irrelevant information, such as axes and scaling maps. Next, all cropped glitches are resized to a standardized size of $448\times 448\times 3$ to facilitate uniformity during model training. Finally, min-max normalization is applied, transforming the pixel intensities into a range of 0 to 1, ensuring that all glitches are equally scaled and enhancing the model’s ability to effectively learn from the data.

### v2

@@ Abstract
The first successful detection of gravitational waves by ground-based observatories, such as the Laser Interferometer Gravitational-Wave Observatory (LIGO), marked a breakthrough in our comprehension of the Universe. However, due to the unprecedented sensitivity required to make such observations, gravitational-wave detectors also capture disruptive noise sources called glitches, which can potentially be confused for or mask gravitational-wave signals. To address this problem, a community-science project, Gravity Spy, incorporates human insight and machine learning to classify glitches in LIGO data. The machine-learning classifier, integrated into the project since 2017, has evolved over time to accommodate increasing numbers of glitch classes. Despite its success, limitations have arisen in the ongoing LIGO fourth observing run (O4) due to the architecture’s simplicity, which led to poor generalization and inability to handle multi-time window inputs effectively. We propose an advanced classifier for O4 glitches. Using data from previous observing runs, we evaluate different fusion strategies for multi-time window inputs, using label smoothing to counter noisy labels, and enhancing interpretability through attention module-generated weights. Our new O4 classifier shows improved performance, and will enhance glitch classification, aiding in the ongoing exploration of gravitational-wave phenomena.
: Class. Quantum Grav.
@@ 1 Introduction
The first discovery of gravitational waves, a pivotal element in Einstein’s theory of general relativity [1], has opened up an entirely new window in the cosmos. The Laser Interferometer Gravitational-Wave Observatory (LIGO) [2] achieved groundbreaking success in detecting these ripples in spacetime for the first time in September 2015 [3]. Since then, the gravitational-wave detector network has been expanded with the addition of Virgo [4], which joined observations in August 2017 [5, 6], and KAGRA [7], which initiated its observing run in April 2020 [8]; this network has collected a large catalog of gravitational-wave observations [9]. However, collecting these observations demands exceptionally sensitive and intricate detectors in order to be able to measure minuscule fluctuations in spacetime [10]. This heightened sensitivity, in turn, leads to the detection of diverse sources of noise, with the potential to obscure or mimic authentic gravitational-wave signals [11, 12, 9, 13]. Bursts of non-Gaussian noise caused by environmental or instrumental factors, known as glitches, are particularly disruptive to measuring gravitational waves. The origins of many glitches remain largely unknown [14, 15, 16]. The process of characterizing and eliminating these glitches is essential to enhancing the quality of the detection system and the analysis of detector data.
@@
In pursuit of a better understanding of detector noise, Gravity Spy [17, 18], a community-science project, has been launched to leverage the power of both volunteers and machine learning to classify glitches in LIGO data. On one front, the project engages human expertise in identifying and categorizing various glitches collected by LIGO. These labeled glitches then serve as training data to refine machine-learning classifiers, enhancing the accuracy and efficiency of glitch identification. Conversely, the classifiers provide guidance to volunteers, assisting them in identifying existing or new potential glitch classes [19]. Gravity Spy is implemented on the Zooniverse platform [20] on the community-science side, and it has demonstrated significant success by attracting over 30 thousand volunteers making more than 7 million glitch classifications.111gravityspy.org The outputs are actively used by LIGO detector-characterization experts [12, 21]. One critical component contributing to the results of Gravity Spy is the integration of the machine-learning classifier.
@@
The machine-learning classifier has been seamlessly integrated into the Gravity Spy pipeline for over five years, largely maintaining its original architecture. Periodically, fine-tuning of its final layer has been performed to enhance its adaptability to diversely predicted glitch classes [22]. The input for the classifier remains unchanged, featuring a structure that juxtaposes four spectrograms of the data from different time windows ($0.5~\mathrm{s}$, $1.0~\mathrm{s}$, $2.0~\mathrm{s}$ and $4.0~\mathrm{s}$, as shown in Figure 1) of a single glitch into one merged image. The same images are shown to volunteers on Zooniverse for their classifications [17].
The first version of the machine-learning classifier was introduced in the initial Gravity Spy paper [17] and was integrated into the Gravity Spy pipeline after the completion of the first observing run (O1). Designed for a 20-class classification setup, this classifier consisted of two convolutional layers and two fully-connected layers to extract image features. Each convolutional layer was followed by a max pooling layer to reduce the dimensionality of the features.
@@
As the project progressed, during the second observing run (O2), Gravity Spy volunteers identified new glitch classes and two of them, 1080 Lines and 1400 Ripples, were added into the classifier. Therefore, the classifier was optimized by incorporating two additional convolutional layers and max pooling layers before the existing fully-connected layers, which facilitated the extraction of more discriminative features and adapted the classifier to a $22$-class classification task [23].
@@
During the third observing run (O3), two new glitch classes, known as Low-frequency Blip and Fast Scattering (or Crown), were identified through the collaborative efforts of volunteers and LIGO detector-characterization experts [24]. Alongside this, the decision was made to remove the None of the Above class from the training dataset. The None of the Above class was initially aimed at empowering our volunteers to identify new glitch classes. However, its inclusion confuses the model, especially given morphological similarities with other glitches. As a result, removing this class from the model allows the identification of new glitches more effectively with low predicted confidence [24, 25]. This led to further retraining of the classifier, enabling it to effectively accommodate a 23-class classification task [24].
@@
Currently, the fourth observing run (O4) of the gravitational-wave detector network is in progress [26, 27, 28]. With the anticipation of discovering new glitch classes during O4 and subsequent runs, the current Gravity Spy classifier faces major limitations [29, 30, 25]. First, in the glitch classification task, unlike other image processing tasks, location plays a crucial role. For example, distinguishing between a line at the top and a line at the bottom of the image is essential as they represent different glitches at different frequencies. However, the classifier’s relatively straightforward architecture, comprised mainly of a combination of convolutional and fully-connected layers, hinders its capacity to capture unique features in individual glitch classes [31]. Second, due to its shallow layer structure, the classifier requires significant resizing of input image dimensions, resulting in the loss of valuable information. Moreover, previous studies [32] have shown that the classification accuracy of glitches varies depending on the time window used. However, earlier classifier studies [17, 32] relied on a single model to extract features from merged image inputs across multiple time windows, limiting their ability to capture cross-time window correlations [33]. Recognizing these challenges, it becomes apparent that the demand to analyze more data and diverse glitch types exceeds the current model’s capabilities.
@@
Additionally, the current classifier grapples with a confidence overfitting issue [34, 35], displaying excessive certainty in glitch predictions even when errors are present. This tendency can potentially mislead both volunteers and LIGO detector-characterization experts, posing a risk to the efficiency and reliability of glitch characterization. Therefore, there is a compelling need for an advanced classifier architecture with robust generalization capabilities in the context of O4.
@@
In this study, we develop a novel machine-learning classifier for glitch classification in the ongoing O4 run to address the challenges faced by Gravity Spy. The proposed classifier employs an attention-based multi-view fusion strategy to capture cross-time window correlations across the four time windows. The incorporation of a regularizer and label smoothing techniques into the loss function improves generalization and mitigates the confidence overfitting issue. The contributions of this paper can be summarized as follows: (1) we comprehensively compare three fusion strategies (early fusion, late fusion, and intermediate fusion) in scenarios with multiple time windows as inputs; (2) we apply label smoothing to mitigate the impact of noisy training labels, and (3) we introduce attention modules in the classifier, enhancing transparency in the glitch decision-making process by identifying specific window time that predominantly influences the glitch classification decision—an exploration undertaken for the first time.
@@ 2 Materials and methods
In this section, we provide an overview of the Gravity Spy dataset, covering its data-generation process, image preprocessing, and the methodology for data splitting during both training and testing phases. Following this, we delve into the multi-view fusion strategy, discussing its early, intermediate, and late fusion components. Last, we propose a novel classifier, outlining the model architecture, attention modules, and the application of a label-smoothing technique.
@@ 2.1 Gravity Spy dataset
The Gravity Spy dataset comprises time–frequency spectrograms of glitches.
The glitches are initially identified by the Omicron pipeline [36]. Omicron searches for excess power in the data by performing a constant-$Q$ transform [37], which projects the data onto a basis of Gaussian-windowed complex-valued exponential functions that tile the time–frequency plane ($Q$ refers to the ratio of duration to bandwidth), and then searching for significant clusters of tiles. Omicron calculates signal-to-noise ratios (SNRs) for triggers from the magnitudes of the most significant time–frequency tiles [38]. To concentrate on glitches that pose significant challenges to gravitational-wave analyses (and are visible in spectrograms), the Gravity Spy pipeline only considers triggers with an Omicron SNR greater than $7.5$, and where the peak frequency falls between $10~\mathrm{Hz}$ and $2048~\mathrm{Hz}$.
@@
Once data of interest have been identified, a constant-$Q$ transform [39, 38] is employed to convert whitened LIGO strain data into time–frequency spectrograms, known as Omega scans [12]. Gravity Spy generates four spectrograms for each glitch triggered, using distinct time windows: $0.5~\mathrm{s}$, $1.0~\mathrm{s}$, $2.0~\mathrm{s}$ and $4.0~\mathrm{s}$. Each window is centred on the Omicron trigger time. The incorporation of multiple time windows serves a crucial purpose, allowing both volunteers and machine-learning models to analyze the morphologies of glitches occurring at different characteristic timescales. For instance, this enables the examination of short- (e.g., Blip) and long-duration (e.g., Whistle) noise transients, as depicted in Figure 1.
@@
Figure 1: Time–frequency spectrogram examples of two glitches with four time windows ($0.5~\mathrm{s}$, $1.0~\mathrm{s}$, $2.0~\mathrm{s}$ and $4.0~\mathrm{s}$): (a) Blip, characterized by its short duration, and (b) Whistle, characterized by its long duration. These examples illustrate how different time windows enable the analysis of glitches occurring at various temporal scales, providing valuable insights into the morphological characteristics of each glitch type.
@@
The training and testing datasets employed in this study are sourced from a set of Gravity Spy project classifications [40], focusing exclusively on data from the Hanford and Livingston detectors during O3. This classification task has in total 23 glitch classes. To ensure consistency, the names associated with each class and the typical morphology of glitches belonging to each class were determined through a collaborative effort between LIGO scientists and citizen scientists. Examples representing each of the 23 glitch classes are illustrated in Figure 2, and more details are provided in the O1–O3 data release paper [25].
@@
Figure 2: Time–frequency morphologies for each of the 23 glitch classes in the Gravity Spy dataset. Each image is a spectrogram, with all sharing the same axes and scale as for the Violin Mode glitch. These examples provide one possible morphological representation for each class, while each class can exhibit various shapes and patterns.
@@
For the purposes of this study, we consider the ground truth for glitches in our dataset to be established through a collaborative effort involving volunteers and the previously implemented machine-learning classifier [17]. However, this ground truth may exhibit a potential noisy-label issue, that is a few glitches were misclassified by the classifier and some volunteers. To strengthen the confidence in the ground truth used for our new classification model, we adopt a filtering criterion to create the training and testing dataset. Specifically, we included only glitches with consensus confidence scores above $0.9$ from both machine-learning models and volunteer classifications, which accounts for $84.5\%$ of the entire dataset. As a consequence, our new model is trained on glitches with relatively reliable labels, effectively reducing confusion during the training process and shielding the model from potentially mislabeled ground truth. Improved training sets for O4 data will be the focus of future studies.
@@
Following the filtering criteria, the Gravity Spy dataset is divided into three distinct sets. The training set consists of $8439$ glitches and the validation set consists of $1687$ glitches. Additionally, a separate set of $3538$ glitches is reserved for testing purposes. Glitches were selected randomly to be partitioned into the three sets. Each set is carefully curated to ensure no overlapping glitch subjects, thereby preventing data leakage between those three phases. A detailed breakdown of the data splitting is presented in Table 1.
@@
Table 1: The distribution of each class in the Gravity Spy O3 dataset across the training, validation, and testing sets.
Dataset size
Gravity Spy class Train Validation Testing
1080 Lines 184 49 225
1400 Ripples 257 64 224
Air Compressor 293 57 225
Blip 312 84 218
Blip Low Frequency 337 103 222
Chirp 38 10 17
Extremely Loud 327 85 224
Fast Scattering 515 128 224
Helix 46 10 71
Koi Fish 338 93 224
Light Modulation 127 28 159
Low-frequency Burst 378 106 223
Low-frequency Lines 381 91 225
No Glitch 311 78 225
Paired Doves 124 31 146
Power Line 354 83 224
Repeating Blips 411 90 224
Scattered Light 521 128 224
Scratchy 178 36 225
Tomte 408 90 225
Violin Mode 210 56 225
Wandering Line 25 8 25
Whistle 462 129 224
@@
The glitches from all three cohorts undergo the same pre-processing steps. Initially, each spectrogram image’s original dimension size is $600\times 800\times 3$ (height, width, color channel). We first crop the image with a bounding box defined as [top-left vertical:60, top-left horizontal:100, height:580, width:675] to retain only the intensity region of interest and eliminate irrelevant information, such as axes and scaling maps. Next, all cropped glitches are resized to a standardized size of $448\times 448\times 3$ using bilinear interpolation to facilitate uniformity during model training. Finally, min–max normalization is applied to transform each pixel intensity $X$ using ${(X-X_{\mathrm{min}})}/{(X_{\mathrm{max}}-X_{\mathrm{min}})}$, where $X_{\mathrm{min}}$ and $X_{\mathrm{max}}$ are the minimal and maximum pixel intensities in the dataset. This transformation scales the pixel intensities to a range of 0 to $1$, ensuring that all images are equally scaled, and enhancing the model’s ability to effectively learn from the data.